# Optimal Trajectory Planning for Cooperative Manipulation with Multiple Quadrotors Using Control Barrier Functions

### v1

@@ V-A Setup
We conducted real-world experiments in a $10\times 6\times 4\leavevmode\nobreak\ $\mathrm{m}$^{3}$ flying space at the ARPL lab at New York University, using a Vicon motion capture system for position feedback of the payload, attach points, and quadrotors at $100\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$. The geometric controller from our previous work [4] controls the system, using the desired payload state from the proposed planning algorithms. The controller distributes cable tension forces to each quadrotor, and the onboard controllers generate motor commands to achieve the required tensions. Fig. 5 illustrates the data flow between the motion capture system, planning algorithm, controller, and quadrotors.
In the real-world experiments, we create a narrow gap using termocol boxes. We use the planner to find a path thorough the gap. We increase the size of the obstacle by $7.5\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}$$ before feeding into the planner to account for uncertainty in the real-world system. On the other hand, in simulator, there was no need to over approximate the shape of the obstacle and we see good obstacle avoidance using our generated trajectory.
@@
Figure 5: System block diagram in the real-world experiments.
@@ V-B Trajectory Generation
We test our algorithm on a triangular payload,which weighted $196\leavevmode\nobreak\ $\mathrm{g}$$, carried by $3$ robots attached to its $3$ corners by inelastic strings. We approximate our payload,strings and quadrotors as convex polytopes. In our experiments, we had a equilateral triangular payload of side 1m, so we took a polytope with top and bottom face same as dimensions of the triangle and a height of $0.1\leavevmode\nobreak\ $\mathrm{m}$$. A plane in form of $\mathbf{A}\mathbf{y}\leq\mathbf{B}$, where $\mathbf{y}\in\mathbb{R}^{3}$ divides the 3D space into two parts. So our polytope which has five planes can be combined together to represent all the points in and on the polytope as $\mathbf{A}\mathbf{y}\leq\mathbf{B}$, where $\mathbf{y}\in\mathbb{R}^{3}$. The cables and cuboid are similarly approximated together as cuboid of dimensions $0.15\leavevmode\nobreak\ $\mathrm{m}$\times 0.15\leavevmode\nobreak\ $\mathrm{m}$\times 1\leavevmode\nobreak\ $\mathrm{m}$$. Similarly, we also model the obstacles in the environment as convex polytopes. We formulate the Optimization problem in Casadi optimizer with horizon length of $25$. We take the first input, apply on the system and then re-solve the optimization problem. Each iteration takes between $9$ and $15\leavevmode\nobreak\ $\mathrm{s}$$ on an Intel(R) Core(TM) i7-8565U CPU. The time depends on the number of obstacles that are active at a given state. We chose to ignore obstacles farther than $0.6\leavevmode\nobreak\ $\mathrm{m}$$ in our planning.
@@ V-C Results
In larger and twisted environments, we observed that our proposed planners often resulted in stucking in locally optimal solutions or reaching states from which recovery was infeasible if we don’t use global planner to provide some sorts of references. When integrating a simple global planner, we mitigated these issues by providing a rough reference path for the system to follow, allowing our proposed algorithms to plan more effectively with $100\%$ success rate.
In the real-world experiments, we observed that the quadrotors manipulate the payload by rotating significantly to navigate through the gap between 2 columns. This maneuver allowed the system to pass through the obstacles. The quadrotors demonstrated coordinated behavior, squeezing through obstacle columns and spreading out after clearing them. These results highlight that the generated trajectory successfully ensures that both the payload and quadrotors avoid collisions throughout the task.
@@ VI Conclusion
This paper presents a comprehensive trajectory planning algorithm for multi-quadrotor systems manipulating a cable-suspended payload in cluttered environments. By incorporating control barrier functions (CBFs) into the planning process, the proposed method ensures obstacle avoidance for the payload, cables, and quadrotors, effectively addressing the challenges posed by the complex dynamics of the system. Our use of convex polytopes to model the system components and the application of the Duality theorem significantly reduce the computational complexity of the optimization process, making it feasible for real world applications. Experimental results in both simulation and real-world environments validate the proposed method’s ability to generate collision-free trajectories.

### v2

@@ V Experimental Results
Figure 5: System block diagram in the real-world experiments.
@@ V-A Setup
We conducted real-world experiments in a $10\times 6\times 4\leavevmode\nobreak\ $\mathrm{m}$^{3}$ flying space at the ARPL lab at New York University, using a Vicon motion capture system for the feedback of the payload, attach points, and quadrotors at $100\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$. The geometric controller from our previous work [4] controls the system, using the desired payload state from the proposed planning algorithms. The controller distributes cable tension forces to each quadrotor, and the onboard controllers generate motor commands to achieve the required tensions. Fig. 5 illustrates the data flow between the motion capture system, planning algorithm, controller, and quadrotors.
In the real-world experiments, we create a narrow gap using termocol boxes. We use the planner to find a path through the gap. We increase the size of the obstacle by $7.5\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}$$ before feeding into the planner to account for uncertainty in the real-world system.
@@ V-B Trajectory Generation
We tested our algorithm on a triangular payload, weighted $196\leavevmode\nobreak\ $\mathrm{g}$$, carried by $3$ robots attached to its $3$ corners via cables. We approximate our payload, cables, and quadrotors as convex polytopes. In our experiments, we had an equilateral triangular payload on side $1$\mathrm{m}$$, so we took a polytope with the top and bottom faces of the same dimensions as the triangle dimensions and a height of $0.1\leavevmode\nobreak\ $\mathrm{m}$$. A plane in the form of $\mathbf{A}\mathbf{y}\leq\mathbf{B}$, where $\mathbf{y}\in\mathbb{R}^{3}$ divides the 3D space into two parts. So, our polytope which has five planes can be combined together to represent all points in and on the polytope as $\mathbf{A}\mathbf{y}\leq\mathbf{B}$, where $\mathbf{y}\in\mathbb{R}^{3}$. The cables and cuboid are similarly approximated together as cuboids of dimensions $0.15\leavevmode\nobreak\ $\mathrm{m}$\times 0.15\leavevmode\nobreak\ $\mathrm{m}$\times 1\leavevmode\nobreak\ $\mathrm{m}$$. Similarly, we also model the obstacles in the environment as convex polytopes. We formulate the optimization problem in the Casadi optimizer with an horizon length of $25$ steps. We take the first input, apply on the system and then re-solve the optimization problem. Each iteration takes between $9$ and $15\leavevmode\nobreak\ $\mathrm{s}$$ on an Intel(R) Core(TM) i7-8565U CPU. The time depends on the number of obstacles that are active in a given state. We chose to ignore obstacles farther than $0.6\leavevmode\nobreak\ $\mathrm{m}$$ in our planning.
@@ V-C Results
In larger and twisted environments, we observed that our proposed planners often stuck in locally optimal solutions or reach states from which recovery was infeasible if we don’t use global planner to provide some sorts of references. When integrating a simple global planner, we mitigated these issues by providing a rough reference path for the system to follow, allowing our proposed algorithms to plan more effectively with a success rate of $100\%$.
In real-world experiments, we observed that the quadrotors manipulate the payload by rotating significantly to navigate through the gap between two columns. This maneuver allowed the system to pass through the obstacles. The quadrotors demonstrated coordinated behavior, squeezing through obstacle columns and spreading out after clearing them. These results highlight that the generated trajectory successfully ensures that both the payload and quadrotors avoid collisions throughout the task.
@@ VI Conclusion
This paper presents a comprehensive trajectory planning algorithm for multi-quadrotor systems manipulating a cable-suspended payload in cluttered environments. By incorporating control barrier functions (CBFs) into the planning process, the proposed method ensures obstacle avoidance for the payload, cables, and quadrotors, effectively addressing the challenges posed by the complex dynamics of the system. We use convex polytopes to model the system components and apply the Duality theorem to reduce the computational complexity of the optimization process, making it feasible for real-world applications. Experimental results in both simulation and real-world environments validate the proposed method’s ability to generate collision-free trajectories.